OpenStudy (anonymous):

3^n = 1/81 What value of "n" solves the equation?

OpenStudy (error1603):

k

OpenStudy (mathmate):

hints: laws of exponents: \(a^{-n}=\frac{1}{a^n}\) this means the negation of the exponent changes the expression to its reciprocal.

OpenStudy (anonymous):

Okay then what @mathmate

OpenStudy (mathmate):

@Jnate621 Do you know how to solve for n in \(3^n=81\)

OpenStudy (anonymous):

thats basically it. like 3^-1 would be 1/3 so 3^-2 would be 1/9 3^-3 would be 1/27. Once you think about it for a minute. its quite easy

OpenStudy (mathmate):

The definition of exponents is" \(a^n = a\times a\times....\times a\) n times. So \(3^n = 3\times 3\times....\times 3\) n times. Can you find n?

OpenStudy (anonymous):

if you have a negative exponent just make it a fraction. Im not quite sure how to explain it but its called law of exponents like math mate showed

OpenStudy (anonymous):

So is N three? Im still confused.

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

is n was three 3^3 would be nine. However -3 is more on the right direction because 3^-3 would b 1/27. ill give you the hint your answer is negative

OpenStudy (mathmate):

Check if n=3 by calculating 3^3=3*3*3

OpenStudy (mathmate):

By the way, IF n=-3, 3^(-3) = 1/(3*3*3)

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

check your answer. what is 3^-4?

OpenStudy (anonymous):

it would basically look like 1/3^4

OpenStudy (anonymous):

1/81?

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

If you say \(3^{-4} = 1/81\), what is the correct answer for n?

OpenStudy (anonymous):

Idk I wanna say its either -4 or 4. I did 3*3*3*3 and got 81.

OpenStudy (mathmate):

The question says: \(3^n = 1/81\) What value of "n" solves the equation? and you say: \(3^{-4} = 1/81\) So what do you think is the answer for n?

OpenStudy (anonymous):